OpenStudy (anonymous):

Solve and check... 2x^2+2x=4

OpenStudy (anonymous):

are you finding the solutions?

OpenStudy (anonymous):

Are you tying to solve for x??

OpenStudy (anonymous):

i mean x=-2 and x=1

OpenStudy (anonymous):

@Jasmine2012 That's wrong

OpenStudy (anonymous):

Solve for x...

OpenStudy (anonymous):

no its right

OpenStudy (anonymous):

What's not right??

OpenStudy (anonymous):

Working with functions....the answers should be a pair of numbers

OpenStudy (anonymous):

yes ther are x=-2,1

OpenStudy (anonymous):

just not sure how to solve the equation

OpenStudy (anonymous):

Well I would do it this way First bring the 4 to the other side of the equation to make a quadratic \[2x^{2}+2x-4=0\] Because the 4 moved to the other side of the equal line it changed signs now solve. What two numbers when multiplied will give you -4 and when added will give you 2??

OpenStudy (anonymous):

2x^2+2x=4 divide both sides by 2, we get x^2 + x = 2 or x^2 + x - 2 = 0 this can be factorised as follows x^2 + 2x - x - 2 = 0 x(x+2) -1(x+2) = 0 (x+2)(x-1)=0 so x= -2 or x= 1

OpenStudy (anonymous):

thats what i said along time ago

OpenStudy (anonymous):

That's wrong

OpenStudy (anonymous):

@Harkirat that's wrong

OpenStudy (anonymous):

2 times -2 = -4 but when added equals 0

OpenStudy (anonymous):

@BABYShawol184 point out what is wrong......

OpenStudy (anonymous):

i got the same thing as Harkirat x=-2 and x=1

OpenStudy (anonymous):

When -2 is put back into the equation the left side and right side don't equal

OpenStudy (anonymous):

NEVERMIND!!! My mistake

OpenStudy (anonymous):

yeah i thought so...

OpenStudy (anonymous):

So sorry made a huge mistake! @Harkirat is right :) Sorry

OpenStudy (anonymous):

2x^2+2x=4 substitute x = -2 LHS = 2*(-2)^2 + 2(-2) = (2*4)-4 = 8-4=4=RHS

OpenStudy (anonymous):

@BABYShawol184 No problem......we all goof up sometimes.... MATHS.....Memory At Times Has (to) Slip...

OpenStudy (anonymous):

Thank you all!!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

@Harkirat Haha :) I'll remember that :)

OpenStudy (anonymous):

yes it was pretty easy but if you need anything else im here :)